# On Discrete Ambiguity Functions of Random Communication Waveforms

### v1

@@ I Introduction
Integrated Sensing and Communication (ISAC) has emerged as a transformative design paradigm for the sixth-generation (6G) wireless networks [challengesfor6G2023, 11206742]. In 2023, the International Telecommunication Union (ITU) released its Recommendation Framework for IMT-2030 and identified ISAC as one of the six primary usage scenarios for 6G [ITU2023]. Unlike conventional systems where sensing and communication are treated as independent functionalities, ISAC envisions a tightly coupled integration of the two to support a wide range of emerging applications, including Smart Cities and Homes, Vehicular Networks, and the Low-Altitude Economy [11206742]. Such integration enables the shared use of frequency and power resources, significantly enhancing spectral and energy efficiency [MengkaitaoTWC, zhang2021overview], while also facilitating mutual performance gains through sensing-assisted communications and communication-assisted sensing [Liufan2023seventyyears].
@@
One of the fundamental challenges in ISAC is to design waveforms that can simultaneously enable accurate target sensing and maintain reliable communication performance. Existing waveform design strategies are generally grouped into three categories: Sensing-centric, communication-centric, and joint design approaches [7855671, Liufan2018optimalwavedesign]. Sensing-centric methods embed communication data into conventional radar signals, while communication-centric schemes repurpose standardized communication waveforms to perform sensing tasks [R.M.M_1963, MR_2003, 8424569, Sturm2011]. Joint design approaches construct novel ISAC waveforms with the objective of achieving a balanced tradeoff between sensing accuracy and communication efficiency [liu-2021-cramer, 9924202, 9724205]. Owing to their inherent compatibility with current cellular protocols, communication-centric designs can be seamlessly integrated into existing network infrastructures without significant hardware modifications, thereby enabling practical ISAC deployment with minimal impact on established communication performance.
@@
In current communication-centric ISAC designs, sensing functionality is largely supported by a small set of physical layer reference signals, such as channel state information reference signals (CSI-RS) and sounding reference signals (SRS) [9805471, 10904093, 10630588]. These signaling elements are deterministic and well structured, which provides desirable auto-and cross-correlation characteristics for delay-Doppler estimation tasks. Nevertheless, they typically account for only around 10% of the entire frame, leaving the majority of resources unused from the sensing perspective. Given the increasingly stringent sensing accuracy requirements foreseen for 6G networks, relying solely on these signals is no longer adequate. It is therefore natural to consider exploiting the remaining 90% time-frequency resources, which primarily transmit communication data payloads, in order to reinforce sensing capability and improve overall resource efficiency. This strategy, however, introduces a fundamental conflict: While deterministic signals are ideal for radar sensing, data payloads are inherently random, and such randomness can significantly degrade sensing performance [LushihangNetwork, LushihangTSP]. This critical issue has recently been identified as the deterministic-random tradeoff (DRT) in ISAC systems [xiong2023fundamental, Xiong2024Magazine]. Understanding how sensing behaves under such randomness is therefore essential. Accordingly, a careful theoretical analysis is required to evaluate the sensing capability when the waveform carries random data.
@@
The ambiguity function (AF), first introduced by Woodward in 1953 [woodward1953probability], characterizes the two-dimensional (2D) self-convolution of a deterministic radar waveform in the delay-Doppler domain. For periodic signals, or signals equipped with a cyclic prefix (CP), the AF can be further generalized to the periodic AF [303781, 144564]. As a foundational tool in radar sensing, the AF enables the evaluation of two key performance metrics: delay-Doppler resolution, governed by the mainlobe width, and mutual interference among targets, determined by the sidelobe level. Consequently, shaping the AF or its zero-Doppler slice (also known as the auto-correlation function (ACF)) to achieve desirable ambiguity characteristics has been a longstanding research topic since the early development of radar systems [11003993, 10935731, 1056980]. Beyond waveform design, the AF is also closely linked to fundamental estimation-theoretic limits. Under a single-target scenario with additive white Gaussian noise (AWGN), maximum likelihood estimation (MLE) reduces to identifying the peak of the AF [kay1998fundamentals1]. Furthermore, the curvature of the AF mainlobe is proportional to the Fisher information and thus directly linked to the Cramér–Rao bound (CRB) [van2004detection]. In addition, the Ziv–Zakai bound (ZZB), widely recognized as a tighter bound than the CRB [556118, 10884827], can be expressed as a nonlinear functional of the AF in the single-target AWGN case [6894212].
@@
In addition to Woodward’s conventional AF definition, the fast-slow-time (FST) ambiguity function (FST-AF) also plays a critical role in sensing performance assessment. Specifically, the fast-time dimension corresponds to the samples within a single radar pulse and is used for delay estimation by measuring their arrival times. The slow-time dimension, in contrast, spans consecutive pulses and enables Doppler estimation through phase variations across pulses [9492001, 10242003, 8793214]. In radar signal processing, the FST representation typically aligns with the widely adopted “stop-and-go” model, where the Doppler shift is assumed constant within each pulse and only varies across pulses [Richards2005Fundamentals]. This approximation is valid in scenarios with small or moderate Doppler frequencies, where inter-carrier interference (ICI) is negligible. The FST-AF is defined as the two-dimensional discrete Fourier transform (2D-DFT) of the element-wise squared signal in the frequency–slow-time domain. At the receiver, the 2D-DFT operation over the echo signal yields the multi-target delay-Doppler profile, which may be interpreted as a noisy linear combination of multiple time-frequency shifted FST-AFs.
@@
In our previous study [liu2024ofdmachieveslowestranging], we derived closed-form expressions for the average squared ACF of random communication signals under both periodic and aperiodic configurations. The results rigorously proved that, among all orthogonal waveforms employing Quadrature Amplitude Modulation (QAM) or Phase Shift Keying (PSK) constellations, OFDM achieves the lowest ranging sidelobe level, thereby enabling superior ranging accuracy. These findings were further extended to pulse-shaped signals in [11037613], where a metaphorical “iceberg hidden in the sea” structure was introduced to describe the statistical behavior of the ACF. Specifically, the “iceberg” corresponds to the squared mean of the ACF, determined by the squared ACF of the adopted pulse-shaping filter, whereas the “sea level”, arising from the randomness of data payloads, represents the variance of the ACF. However, these studies were restricted to the zero-Doppler characteristics of the AF. A complete characterization of the ambiguity properties of random communication signals in the full delay-Doppler domain therefore remains largely unexplored. This necessitates a deeper investigation into the statistical ambiguity properties of communication-centric ISAC waveforms, beyond conventional deterministic waveform analysis for radar systems.
@@ II-B2 DP-AF
where the expectation is taken over the random symbol vector $\mathbf{s}$. In sight of $|\mathcal{A}_{\rm{DP}}(0,0)|^{2}=\|\mathbf{x}\|_{2}^{4}$, the average mainlobe level may be computed by [liu2024ofdmachieveslowestranging]
@@ III-A Main Results
It is observed that the volume of DP-AF is $N$ times that of its mainlobe, which directly follows from the invariance of the volume of Woodward’s AF [woodward1953probability]. Then we have
@@ III-B1 Average Squared DP-AF for OFDM
As indicated in the previous discussion, for sub-Gaussian constellations ($1\leq\kappa<2$), the N1N-1 lowest sidelobes are all placed in the zero-Doppler slice of the DP-AF of OFDM. This is consistent with the results in [liu2024ofdmachieveslowestranging], that CP-OFDM achieves the lowest average ranging sidelobe among all orthogonal waveforms with CP.
@@ IV-C Optimal Signaling Schemes under FST-AF Regime
where the maximum is attained if $\mathbf{V}^{T}\otimes\mathbf{F}_{N}\mathbf{U}=\mathbf{I}_{NM}$ [liu2024ofdmachieveslowestranging]. This can be realized by letting
@@
where the minimum is attained if all the entries of $\mathbf{V}^{T}\otimes\mathbf{F}_{N}\mathbf{U}$ have constant modulus [liu2024ofdmachieveslowestranging]. In such a case, an achieving strategy would be
@@
Remark 6. We note that the above optimality of OFDM and OTFS can be regarded as a two-dimensional extension of the findings in [liu2024ofdmachieveslowestranging], which show that OFDM and SC yield the lowest average ranging sidelobe levels under sub-Gaussian and super-Gaussian constellations, respectively. The analogy stems from the fact that, within the FST-AF framework, OTFS can be seen as the Fourier dual of OFDM under the 2D signaling schemes, just as SC is the 1D Fourier dual of OFDM.
@@ Appendix A Proof of Proposition 2
By using [liu2024ofdmachieveslowestranging, Proposition 1], $\mathbb{E}\left(\tilde{\mathbf{s}}{\tilde{\mathbf{s}}}^{H}\right)$ may be decomposed as
@@ Appendix G Proof of Proposition 7
where $\tilde{\mathbf{S}}=\mathbb{E}\left\{\operatorname{vec}\left({\mathbf{s}}{\mathbf{s}}^{H}\right)\operatorname{vec}^{H}\left({\mathbf{s}}{\mathbf{s}}^{H}\right)\right\}$, which can be again decomposed as [liu2024ofdmachieveslowestranging]

### v2

@@ I Introduction
Integrated Sensing and Communication (ISAC) has emerged as a transformative design paradigm for the sixth-generation (6G) wireless networks [1, 2]. In 2023, the International Telecommunication Union (ITU) released its Recommendation Framework for IMT-2030 and identified ISAC as one of the six primary usage scenarios for 6G [3]. Unlike conventional systems where sensing and communication are treated as independent functionalities, ISAC envisions a tightly coupled integration of the two to support a wide range of emerging applications, including Smart Cities and Homes, Vehicular Networks, and the Low-Altitude Economy [2]. Such integration enables the shared use of frequency and power resources, significantly enhancing spectral and energy efficiency [4, 5], while also facilitating mutual performance gains through sensing-assisted communications and communication-assisted sensing [6].
@@
One of the fundamental challenges in ISAC is to design waveforms that can simultaneously enable accurate target sensing and maintain reliable communication performance. Existing waveform design strategies are generally grouped into three categories: Sensing-centric, communication-centric, and joint design approaches [7, 8]. Sensing-centric methods embed communication data into conventional radar signals, while communication-centric schemes repurpose standardized communication waveforms to perform sensing tasks [9, 10, 11, 12]. Joint design approaches construct novel ISAC waveforms with the objective of achieving a balanced tradeoff between sensing accuracy and communication efficiency [13, 14, 15]. Owing to their inherent compatibility with current cellular protocols, communication-centric designs can be seamlessly integrated into existing network infrastructures without significant hardware modifications, thereby enabling practical ISAC deployment with minimal impact on established communication performance.
@@
In current communication-centric ISAC designs, sensing functionality is largely supported by a small set of physical layer reference signals, such as channel state information reference signals (CSI-RS) and sounding reference signals (SRS) [16, 17, 18]. These signaling elements are deterministic and well structured, which provides desirable auto-and cross-correlation characteristics for delay-Doppler estimation tasks. Nevertheless, they typically account for only around 10% of the entire frame, leaving the majority of resources unused from the sensing perspective. Given the increasingly stringent sensing accuracy requirements foreseen for 6G networks, relying solely on these signals is no longer adequate. It is therefore natural to consider exploiting the remaining 90% time-frequency resources, which primarily transmit communication data payloads, in order to reinforce sensing capability and improve overall resource efficiency. This strategy, however, introduces a fundamental conflict: While deterministic signals are ideal for radar sensing, data payloads are inherently random, and such randomness can significantly degrade sensing performance [19, 20]. This critical issue has recently been identified as the deterministic-random tradeoff (DRT) in ISAC systems [21, 22]. Understanding how sensing behaves under such randomness is therefore essential. Accordingly, a careful theoretical analysis is required to evaluate the sensing capability when the waveform carries random data.
@@
The ambiguity function (AF), first introduced by Woodward in 1953 [23], characterizes the two-dimensional (2D) self-convolution of a deterministic radar waveform in the delay-Doppler domain. For periodic signals, or signals equipped with a cyclic prefix (CP), the AF can be further generalized to the periodic AF [24, 25]. As a foundational tool in radar sensing, the AF enables the evaluation of two key performance metrics: delay-Doppler resolution, governed by the mainlobe width, and mutual interference among targets, determined by the sidelobe level. Consequently, shaping the AF or its zero-Doppler slice (also known as the auto-correlation function (ACF)) to achieve desirable ambiguity characteristics has been a longstanding research topic since the early development of radar systems [26, 27, 28]. Beyond waveform design, the AF is also closely linked to fundamental estimation-theoretic limits. Under a single-target scenario with additive white Gaussian noise (AWGN), maximum likelihood estimation (MLE) reduces to identifying the peak of the AF [29]. Furthermore, the curvature of the AF mainlobe is proportional to the Fisher information and thus directly linked to the Cramér–Rao bound (CRB) [30]. In addition, the Ziv–Zakai bound (ZZB), widely recognized as a tighter bound than the CRB [31, 32], can be expressed as a nonlinear functional of the AF in the single-target AWGN case [33].
@@
In addition to Woodward’s conventional AF definition, the fast-slow-time (FST) ambiguity function (FST-AF) also plays a critical role in sensing performance assessment. Specifically, the fast-time dimension corresponds to the samples within a single radar pulse and is used for delay estimation by measuring their arrival times. The slow-time dimension, in contrast, spans consecutive pulses and enables Doppler estimation through phase variations across pulses [34, 35, 36]. In radar signal processing, the FST representation typically aligns with the widely adopted “stop-and-go” model, where the Doppler shift is assumed constant within each pulse and only varies across pulses [37]. This approximation is valid in scenarios with small or moderate Doppler frequencies, where inter-carrier interference (ICI) is negligible. The FST-AF is defined as the two-dimensional discrete Fourier transform (2D-DFT) of the element-wise squared signal in the frequency–slow-time domain. At the receiver, the 2D-DFT operation over the echo signal yields the multi-target delay-Doppler profile, which may be interpreted as a noisy linear combination of multiple time-frequency shifted FST-AFs.
@@
In our previous study [38], we derived closed-form expressions for the average squared ACF of random communication signals under both periodic and aperiodic configurations. The results rigorously proved that, among all orthogonal waveforms employing Quadrature Amplitude Modulation (QAM) or Phase Shift Keying (PSK) constellations, OFDM achieves the lowest ranging sidelobe level, thereby enabling superior ranging accuracy. These findings were further extended to pulse-shaped signals in [39], where a metaphorical “iceberg hidden in the sea” structure was introduced to describe the statistical behavior of the ACF. Specifically, the “iceberg” corresponds to the squared mean of the ACF, determined by the squared ACF of the adopted pulse-shaping filter, whereas the “sea level”, arising from the randomness of data payloads, represents the variance of the ACF. However, these studies were restricted to the zero-Doppler characteristics of the AF. A complete characterization of the ambiguity properties of random communication signals in the full delay-Doppler domain therefore remains largely unexplored. This necessitates a deeper investigation into the statistical ambiguity properties of communication-centric ISAC waveforms, beyond conventional deterministic waveform analysis for radar systems.
@@ II-B2 DP-AF
where the expectation is taken over the random symbol vector $\mathbf{s}$. In sight of $|\mathcal{A}_{\rm{DP}}(0,0)|^{2}=\|\mathbf{x}\|_{2}^{4}$, the average mainlobe level may be computed by [38]
@@ III-A Main Results
It is observed that the volume of DP-AF is $N$ times that of its mainlobe, which directly follows from the invariance of the volume of Woodward’s AF [23]. Then we have
@@ III-B1 Average Squared DP-AF for OFDM
As indicated in the previous discussion, for sub-Gaussian constellations ($1\leq\kappa<2$), the N1N-1 lowest sidelobes are all placed in the zero-Doppler slice of the DP-AF of OFDM. This is consistent with the results in [38], that CP-OFDM achieves the lowest average ranging sidelobe among all orthogonal waveforms with CP.
@@ IV-C Optimal Signaling Schemes under FST-AF Regime
where the maximum is attained if $\mathbf{V}^{T}\otimes\mathbf{F}_{N}\mathbf{U}=\mathbf{I}_{NM}$ [38]. This can be realized by letting
@@
where the minimum is attained if all the entries of $\mathbf{V}^{T}\otimes\mathbf{F}_{N}\mathbf{U}$ have constant modulus [38]. In such a case, an achieving strategy would be
@@
Remark 6. We note that the above optimality of OFDM and OTFS can be regarded as a two-dimensional extension of the findings in [38], which show that OFDM and SC yield the lowest average ranging sidelobe levels under sub-Gaussian and super-Gaussian constellations, respectively. The analogy stems from the fact that, within the FST-AF framework, OTFS can be seen as the Fourier dual of OFDM under the 2D signaling schemes, just as SC is the 1D Fourier dual of OFDM.
@@ Appendix A Proof of Proposition 2
By using [38, Proposition 1], $\mathbb{E}\left(\tilde{\mathbf{s}}{\tilde{\mathbf{s}}}^{H}\right)$ may be decomposed as
@@ Appendix G Proof of Proposition 7
where $\tilde{\mathbf{S}}=\mathbb{E}\left\{\operatorname{vec}\left({\mathbf{s}}{\mathbf{s}}^{H}\right)\operatorname{vec}^{H}\left({\mathbf{s}}{\mathbf{s}}^{H}\right)\right\}$, which can be again decomposed as [38]